OpenStudy (iwatermelon):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. x + (y + 5) = 6 B. (y + 5) + y = 6 C. x = (x – 6) + 5 D. x + y = y + 5

alones (alones):

x + y = 6 x = y + 5 let's lok at the first one x+y=6 Step 1: Add -y to both sides. x+y+−y=6+−y x=−y+6 turns in to x + y = 6 also substitute this one x = y + 5 and you will get (y + 5) + y =6 Put them together so it would look like this (y + 5) + y =6